The Coffee Club sells specialist coffee drinks from a rented cart on the beachside on the Gold Coast. Provided below is a summarised version of the income statement for July 2016

Sales revenue

$7,500

Less: Expenses

Cost of beverages

$750

Cost of napkins

750

Rent of cart

750

Wages

375

Total expenses

2,625

Profit

$4,875

Collapse question part

(a) Calculate the contribution margin ratio. (Round answer to the nearest whole number.)

Contribution margin ratio %

(b) calculate the breakeven sales in dollars ?
Answer #1

There should be a segregation of variable costs and fixed cost.

Variable costs are directly varied with production units; therefore, these are cost of beverages, cost of napkins, and wages.

Fixed cost is the amount of rent, since it doesn’t vary with production units.

Total contribution = Sales revenues – Total variable cost

                                    = 7,500 – (750 + 750 + 375)

                                    = 7,500 – 1,875

                                    = 5,625

Contribution margin ratio = (Total contribution / Sales revenue) × 100

                                          = (5,625 / 7,500) × 100

                                          = 75% (Answer)

Breakeven sales = Fixed cost / Contribution margin ratio

                           = 750 / 75%

                           = 750 / 0.75

                           = $1,000 (Answer)